A firm manufactures two products; the net profit on product 1 is Birr 3 per unit and Birr 5 per unit on product 2. The manufacturing process is such that each product has to be processed in two departments D1 and D2. Each unit of product1 requires processing for 1 minute at D1 and 3 minutes at D2; each unit of product 2 requires processing for 2 minutes at D1 and 2 minutes at D2. Machine time available per day is 860 minutes at D1 and 1200 minutes at D2. How much of product 1 and 2 should be produced every day so that total profit is maximum. (solve with graphical method)


1
Expert's answer
2022-02-03T17:26:22-0500

Solution:

Let "x_1" and "x_2" be levels of production of two products, then

"Z=3 x_1+5 x_2" is profit function and it should be maximized.

Subject to the constraints:

"x_1+2 x_2\le860;\\
3 x_1+2 x_2\le1200;\\"

"x_1\ge0,x_2\ge0"


Corners points are (0,430), (170,345), (400,0).

At (0,430), "Z=3(0)+5 (430)=2150"

At (170,345), "Z=3(170)+5 (345)=2235"(Maximum)

At (400,0), "Z=3(400)+5 (0)=1200"

Hence, the optimal solution to the given LP problem is : x1=170,x2=345 and max Z=2235.